HIGH-RESOLUTION INFRARED SPECTROSCOPY OF THE (1, 0, 1) − (0, 0, 0) BAND OF C_{3}

\underline{S. THORWIRTH}, J. KRIEG, I. KEPPELER, V. LUTTER, S. SCHLEMMER, T. F. GIESEN, I. Physikalisches Institut, Universität zu Köln, 50937 Köln, Germany; M. E. HARDING, Karlsruher Institut für Technologie, Institut für Nanotechnologie, 76021 Karlsruhe, Germany; J. VÁZQUEZ, Center for Theoretical Chemistry, Department of Chemistry and Biochemistry, The University of Texas at Austin, Austin, Texas 78712, U.S.A.

Using the carbon cluster experiment at Cologne in combination with a home-made optical parametric oscillator operating at three microns, the (1, 0, 1) − (0, 0, 0) combination band of the C_{3} cluster has been investigated around 3260 cm^{−1}. In addition, the associated (1, 1, 1) − (0, 1, 0) hot band has been observed. Using sample rods enriched in ^{13}C, the (1, 0, 1)−(0, 0, 0) band was also studied for the two isotopologs^{13}CCC and C^{13}CC. The experimental molecular parameters obtained compare very favorably with results from coupled-cluster calculations.